# Discord-type quantum correlations in axially symmetric spin-(1/2, $S$) systems

### v1

@@ VI.3 Sadden change phenomena
Finally, let’s consider a model with slightly different parameters, namely, $B_{1}$, $B_{2}$, $J$, $J_{z}$, $K_{1}$ and $D_{z}$ remain the same, but $K_{2}$ is replaced by -0.6. Look at Fig. 6. Here, each pair of 0- and 1-branches intersects at two points. As a result, when the temperature changes from zero to infinity, the quantum correlations $\cal U$ and $\cal F$ undergo two sudden transitions. Figure 6 clearly shows the intersections of the curves at low temperatures. The $\cal U$ branches (blue curves) intersect at temperatures 0.164 and 0.06 for $S$=1 and 3/2, respectively. Similarly for $\cal F$-branches (orange curves): they intersect at temperatures 0.375 and 0.13 for spin lengths $S$=1 and 3/2, respectively. Unfortunately, at higher temperatures the intersections become less visible, so their surroundings are shown in the insets. From these insets it is clear that at high temperatures the behavior of quantum correlations is determined by 0-branches. The same conclusion follows from the high-temperature decompositions (49) - (50).
VII Summary and outlook
In this paper, we have derived analytical formulas (24), (25), (35) and (IV.2) for the branches of quantum correlations LQU and LQFI. The obtained formulas are surprisingly compact and allow one to calculate the local quantum uncertainty and local quantum Fisher information for any spin-(1/2,$S$) systems whose states commute with the $z$-component of the total spin.
The derived formulas were then applied to the systems in thermal equilibrium, for which the behavior of quantum correlations has been studied in detail. Among findings, we note the following. It is natural to expect that an increase in the spin length, $S\to\infty$, should destroy quantumness of the system. Our calculations confirm this observation, which is known for antiferromagnets at ground states. However, contrary to intuitive expectations, we found enhancement of quantum correlations at zero temperature for systems with ferromagnetic couplings.
We also carefully studied the phenomena of sudden transitions where smooth variations in temperature lead to sharp changes in the behavior of quantum correlations. Remarkably, both single such transitions and series of them have been revealed.
@@ VII Summary and outlook
The obtained formulas can be used to study other properties of quantum correlations, such as their dynamical behavior.
Acknowledgments: S. H. was supported by Semnan University under Contract No. 21270. M.A.Y. was supported in part by a state task, the state registration number of the Russian Federation is ##\##124013000760-0.
CRediT authorship contribution statement
@@
Saeed Haddadi: Methodology, Investigation. M. A. Yurischev: Writing - original draft, Conceptualization.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data was used for the research described in the article.

### v2

@@ V.3 Sadden change phenomena
Finally, let us consider a model with slightly different parameters, namely, $B_{1}$, $B_{2}$, $J$, $J_{z}$, $K$, $K_{1}$, $D_{z}$, $\Gamma$ and $\Lambda$ remain the same, but $K_{2}$ is replaced by $-0.6$. Look at Fig. 5. Here, each pair of 0- and 1-branches intersects at two points. As a result, when the temperature changes from zero to high values, the quantum correlations $\cal U$ and $\cal F$ undergo two sudden transitions. Figure 5 clearly shows the intersections of the curves at low temperatures. The $\cal U$ branches (blue curves) intersect at temperatures 0.164 and 0.06 for $S$=1 and 3/2, respectively. Similarly for $\cal F$-branches (orange curves): they intersect at temperatures 0.375 and 0.13 for spin lengths $S$=1 and 3/2, respectively. Unfortunately, at higher temperatures the intersections become less visible, so their surroundings are shown in the insets. From these insets it is clear that at high temperatures the behavior of quantum correlations is determined by 0-branches. The same conclusion follows from the high-temperature decompositions (34) - (37).
In conclusion of this subsection, we would like to note the following. Although the LQU and LQFI are discord-type quantum correlations, they are determined only by two optimal measurement angles, namely zero and $\pi/2$. Meanwhile, the original (entropic) discord may have an optimal measurement angle between zero and $\pi/2$. This leads to sudden transitions, which are described by Landau’s theory of phase transitions of the first and second kind [64].
VI Summary and outlook
In this paper, we have derived analytical formulas (13), (14), (24) and (25) for the branches of quantum correlations LQU and LQFI. These formulas turned out to be unexpectedly compact, as a result of which computer code for them is written very simply. This actually makes it possible to study the local quantum uncertainty $\cal U$ and the local quantum Fisher information $\cal F$ for any spin-$(1/2,S)$ systems in axisymmetric states. The state for a given $S$ contains $18S+$ 1 free parameters. This allows the investigation of many physically important models.
The derived formulas were then applied to the systems in thermal equilibrium, for which the behavior of quantum correlations depending on temperature was studied. The high-temperature asymptotic behavior of quantum correlations was established, whence it follows that both LQU and LQFI decrease monotonically with increasing $T$ according to the power law $T^{-2}$ and completely disappear at an infinitely high temperature. On the other hand, quantum correlations can increase in finite (limited) regions with increasing temperature. Additionally, we showed that temperature can induce quantum correlations from an uncorrelated ground state.
Next, we studied the effects of sudden changes in quantum correlations for different values of $S$. The performed calculations showed that both LQU and LQFI can experience one or more such transitions.
In conclusion, we investigated the quantum correlations LQU and LQFI and found that they allow us to track the persistence of quantum correlations at finite temperatures and reveal intricate thermal effects such as correlation enhancement and abrupt transitions. However, it is not possible to state precisely which measure is more efficient for quantifying quantum correlations as it depends on the model and parameters of the system under consideration.
@@ VI Summary and outlook
The obtained formulas can be used to study other properties of quantum correlations, for example, their dynamical behavior. We believe that axially symmetric quantum states of qubit-qudit systems will be as useful as X states for two-qubit models.
Acknowledgements
E.I.K. and M.A.Y. were supported in part by a state task, the state registration number of the Russian Federation is ##\##124013000760-0. S.H. thanks the School of Particles and Accelerators at the Institute for Research in Fundamental Sciences for their financial support.
Data availability statement
All data that support the findings of this study are included within the article.
Author contributions
@@ Acknowledgements
Saeed Haddadi: Methodology, Investigation, Writing – review & editing. Elena I. Kuznetsova: Calculations, Interpretation of results. M. A. Yurischev: Writing - original draft, Conceptualization. Thorough checking of the manuscript was done by all authors.
Conflict of interest
The authors declare that no conflicts of interest or personal relationships have influenced this work.
@@
Appendix A Diagonalization of the density matrix
Since the density matrix $\rho$ commutes with the diagonal matrix ${\cal S}_{z}$, it has nonzero matrix elements only between states with the same eigenvalue of ${\cal S}_{z}$. Here, ${\cal S}_{z}$ plays a role of “constant or integral of motion”; in this connection see, e.g., Chapter 1 in the books [65, 66]. As mentioned in Sec. II, ${\cal S}_{z}$ has two non-degenerate and $4S$ doubly degenerate eigenvalues. Therefore, the density matrix (2) actually consists of two one-by-one and $2S$ two-by-two subblocks.444 This will be obvious if to reorder the vectors of the original basis so that the $k$th and $(2S+k)$th vectors become neighbors, then the density matrix (2) takes an explicit quasidiagonal form $\hat{\rho}=[p_{0},\left(\begin{array}[]{cc}a_{1}&u_{1}\\ u_{1}^{*}&a_{2S+1}\\ \end{array}\right),\ldots,\left(\begin{array}[]{cc}a_{2S}&u_{2S}\\ u_{2S}^{*}&a_{4S}\\ \end{array}\right),p_{4S+1}].$ This form is implemented for the system $(S,1/2)$. The first and $(4S+1)$th “subblocks” have sizes $1\times 1$, i.e. they are ready eigenvalues. The remaining $2S$ subblocks are $2\times 2$-dimensional, and their diagonalization is easily accomplished by directly calculating the eigenvectors of the corresponding subblocks (see, for example, [67]). Eigenvectors are constructed using the $k$th and $(2S+k)$th pair of vectors of the original basis. Indeed, a suitable fragment extracted from the inner $4S\times 4S$ block of the density matrix (2) reads
and then the corresponding eigenvectors are given as
where
where ${\tilde{q}}_{k}$ and ${\tilde{u}}_{k}$ ($k=1,\ldots,2S$) are defined by Eq. (44). Note that $R$ has the axial symmetrical structure. Moreover, it is clear that $R^{\dagger}=R$ and $R^{\dagger}R=I$, so this diagonalizing transformation is both Hermitian and unitary.
Using Eqs. (2) and (45) we find
where $p_{1},\ldots,p_{4S}$ are the eigenvalues given by Eq. (2). From here and Eq. (44), it follows that
$q_{k}=p_{k}-a_{2S+k}$ (47)
and
where as before $k=1,\ldots,2S$.
Appendix B Spin matrices in the diagonal representation of the density matrix
Let us transform the local Pauli matrices $\sigma_{\mu}\otimes I_{2S+1}$ ($\mu=x,y,z$) of the particle $A$ into a diagonal representation of the density operator. Using Eq. (45), we give these matrices in open form
and
Unlike the first two matrices, the latter has the axial symmetrical form.